### v2

# Extensions of semibricks and maximal finite semibricks in quiver representations

Abstract.
A semibrick is a set of modules satisfying Schur’s Lemma. We prove that any finite semibrick $\mathcal{S}$ in representations of a finite acyclic quiver can be extended to an infinite semibrick if there exists a brick in $\mathcal{S}$ which is not exceptional. Therefore any maximal finite semibrick consists of exceptional modules.
1. Our results
In this short paper, we discuss extensions of finite semibricks in representations of a finite acyclic quiver $Q$, or equivalently, finite semibricks in the module category $\operatorname{\mathsf{mod}}KQ$ of the path algebra $KQ$. For a finite dimensional algebra $\varLambda$ over an algebraically closed field $K$, bricks, semibricks and related notions are defined as follows:
  • A $\varLambda$-module $B\in\operatorname{\mathsf{mod}}\varLambda$ is called a brick if the endomorphism ring $\operatorname{\mathsf{End}}_{\varLambda}(B)$ is isomorphic to $K$. We write $\operatorname{\mathsf{brick}}\varLambda$ for the set of isoclasses of bricks in $\operatorname{\mathsf{mod}}\varLambda$.
  • [C1] A module $B\in\operatorname{\mathsf{mod}}\varLambda$ is said to be exceptional if $B\in\operatorname{\mathsf{brick}}\varLambda$ and $\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(B,B)=0$.
  • [A] A subset $\mathcal{S}\subset\operatorname{\mathsf{brick}}\varLambda$ is called a semibrick if $\operatorname{\mathsf{Hom}}_{\varLambda}(B,B^{\prime})=0$ for any $B\neq B^{\prime}\in\mathcal{S}$. We write $\operatorname{\mathsf{sbrick}}\varLambda$ for the set of semibricks in $\operatorname{\mathsf{mod}}\varLambda$.
  • A semibrick $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ is said to be finite if $\mathcal{S}$ is a finite set.
  • A semibrick $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ is said to be maximal if there exists no semibrick $\mathcal{S}^{\prime}$ such that $\mathcal{S}\subsetneq\mathcal{S}^{\prime}$.
The following are the main results of this paper.
Theorem 1.1.
Let $\varLambda=KQ$ with $Q$ a finite acyclic quiver, and $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ be a finite semibrick. Assume that a brick $B\in\mathcal{S}$ is not exceptional. Then the following statements hold.
  1. (1)
    There exists some $B^{\prime}\in\operatorname{\mathsf{brick}}\varLambda\setminus\{B\}$ satisfying
    $\mathcal{S}\sqcup\{B^{\prime}\}\in\operatorname{\mathsf{sbrick}}\varLambda,\quad\operatorname{\underline{\mathsf{dim}}}B^{\prime}=\operatorname{\underline{\mathsf{dim}}}B.$
  2. (2)
    There exists an infinite semibrick $\mathcal{S}^{\prime}\in\operatorname{\mathsf{sbrick}}\varLambda$ such that
    $\mathcal{S}\subset\mathcal{S}^{\prime},\quad\text{$\operatorname{\underline{\mathsf{dim}}}B^{\prime}=\operatorname{\underline{\mathsf{dim}}}B$ for each $B^{\prime}\in\mathcal{S}^{\prime}\setminus\mathcal{S}$}.$
Corollary 1.2.
Let $\varLambda=KQ$ with $Q$ a finite acyclic quiver, and $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ be a maximal finite semibrick. Then every brick in $\mathcal{S}$ is exceptional.
For any module $M\in\operatorname{\mathsf{mod}}\varLambda$, we set two subcategories
${{}^{\perp}M}:=\{X\in\operatorname{\mathsf{mod}}\varLambda\mid\operatorname{\mathsf{Hom}}_{\varLambda}(X,M)=0\},\quad M^{\perp}:=\{X\in\operatorname{\mathsf{mod}}\varLambda\mid\operatorname{\mathsf{Hom}}_{\varLambda}(M,X)=0\}.$
Theorem 1.1 and Corollary 1.2 are direct consequences of the following general result.
Theorem 1.3.
Let $\varLambda=KQ$ with $Q$ a finite acyclic quiver, and $M,N\in\operatorname{\mathsf{mod}}\varLambda$. Assume that $B\in{{}^{\perp}M}\cap N^{\perp}$ and that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional. Then there exists some $B^{\prime}\in\operatorname{\mathsf{brick}}\varLambda$ such that
$B^{\prime}\in{{}^{\perp}(M\oplus B)}\cap(N\oplus B)^{\perp},\quad\operatorname{\underline{\mathsf{dim}}}B^{\prime}=\operatorname{\underline{\mathsf{dim}}}B.$
The next section is devoted to showing this. Our key ingredient is [S, Theorem 6.1].
In Section 3, we prove a weaker version of Theorem 1.3 (Theorem 3.1) by using semi-invariants as a record, which was our first proof of Corollary 1.2.
2. Extensions of semibricks and maximal finite semibricks in quiver representations
2.1. Preliminaries
Throughout this paper, let $Q=(Q_{0},Q_{1})$ be a finite acyclic quiver, and $\varLambda=KQ$ be its path algebra over an algebraically closed field $K$. For any module $M\in\operatorname{\mathsf{mod}}\varLambda$, the element $\operatorname{\underline{\mathsf{dim}}}M:=(\operatorname{\mathsf{dim}}_{K}Me_{i})_{i\in Q_{0}}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ is called the dimension vector of $M$, where $e_{i}\in KQ$ is the idempotent corresponding to $i\in Q_{0}$.
Definition 2.1.
Let $d=(d_{i})_{i\in Q_{0}}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then we define the representation variety $\operatorname{\mathsf{rep}}(Q,d)$ of the quiver $Q$ for $d$ as the direct product
$\operatorname{\mathsf{rep}}(Q,d):=\prod_{\alpha\in Q_{1}}\operatorname{\mathsf{Hom}}_{K}(K^{d_{s(\alpha)}},K^{d_{t(\alpha)}}),$
where $s(\alpha)$ and $t(\alpha)$ are the source and the target of each arrow $\alpha\in Q_{1}$. This is an irreducible algebraic variety equipped with the Zariski topology.
In this paper, we will frequently use the following basic fact of $\operatorname{\mathsf{rep}}(Q,d)$, which is immediate from upper semicontinuity.
Lemma 2.2.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$, and $M\in\operatorname{\mathsf{mod}}\varLambda$. Then the following sets are open in $\operatorname{\mathsf{rep}}(Q,d)$:
$\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ :=$:=\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid X\in\operatorname{\mathsf{brick}}\varLambda\},$
$\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}M}$ :=$:=\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid X\in{{}^{\perp}M}\},$
$\operatorname{\mathsf{rep}}(Q,d)\cap M^{\perp}$ :=$:=\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid X\in{M^{\perp}}\}.$
Under this notation, we define Schur roots.
Definition 2.3.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then we call $d$ a Schur root if $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ is nonempty.
By Lemma 2.2, if $d$ is a Schur root, then $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)$.
Schur roots depend on the orientation of the quiver $Q$, but do not depend on the field $K$. This follows from [S, Proposition 6.1] (Proposition 2.5 of this paper) together with the correction [C2] of [S, Proposition 3.3].
We use the Grothendieck group $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$.
For each vertex $i\in Q_{0}$, we write $S_{i}$ for the simple $\varLambda$-module corresponding to $i$, and $[S_{i}]$ for its isoclass. Then the Grothendieck group $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ is the free abelian group $\bigoplus_{i\in Q_{0}}\mathbb{Z}[S_{i}]$, so we get $K_{0}(\operatorname{\mathsf{mod}}\varLambda)\simeq\mathbb{Z}^{Q_{0}}$ as abelian groups. Under this isomorphism, for each module $M\in\operatorname{\mathsf{mod}}\varLambda$, the element $[M]\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ corresponds to its dimension vector $\operatorname{\underline{\mathsf{dim}}}M\in\mathbb{Z}^{Q_{0}}$. Thus in what follows, the dimension vector $\operatorname{\underline{\mathsf{dim}}}M$ is regarded as an element of $K_{0}(\operatorname{\mathsf{mod}}\varLambda)_{\geq 0}:=\sum_{i\in Q_{0}}\mathbb{Z}_{\geq 0}[S_{i}]$.
We set $K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}:=\operatorname{\mathsf{Hom}}_{\mathbb{Z}}(K_{0}(\operatorname{\mathsf{mod}}\varLambda),\mathbb{Z})$. Then each $\theta\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}$ can be considered as a weight, and give the following condition for modules.
Definition 2.4.
[Ki, Definition 1.1] Let $\theta\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}$. For each $X\in\operatorname{\mathsf{mod}}\varLambda$, we say that $X$ is $\theta$-stable if $X\neq 0$, $\theta(X)=0$ and $\theta(Y)>0$ for every quotient module $Y\neq 0,X$ of $X$.
For any $\theta\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}$ and $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$, the set
[…]
is open in $\operatorname{\mathsf{rep}}(Q,d)$ by [S, Lemma 3.1]. On this set, we have the following result.
Proposition 2.5.
[…]
Proof.
The original proof of [S, Theorem 3.3] works in the case that the characteristic of $K$ is zero, but it still holds for any arbitrary algebraically closed field thanks to Crawley-Boevey’s correction [C2]. Then the proof of [S, Theorem 6.1] is valid. ∎
We also use the orbit $\mathcal{O}_{M}$ of each $M\in\operatorname{\mathsf{rep}}(Q,d)$ defined by
$\mathcal{O}_{M}:=\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid X\simeq M\}.$
Since $\varLambda=KQ$ is a hereditary algebra, we have the following property by [Gab, V].
Lemma 2.6 (Voigt’s Lemma).
Let $M\in\operatorname{\mathsf{mod}}\varLambda$ satisfy $\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(M,M)\neq 0$. Then $\mathcal{O}_{M}$ is not dense in $\operatorname{\mathsf{rep}}(Q,d)$, and $\operatorname{\mathsf{rep}}(Q,d)\setminus\overline{\mathcal{O}_{M}}$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)$.
2.2. Proof of Theorem 1.3
To prove Theorem 1.3, we recall some fundamental results on torsion pairs. For each $X\in\operatorname{\mathsf{mod}}\varLambda$, we define $\operatorname{\mathsf{T}}(X)\subset\operatorname{\mathsf{mod}}\varLambda$ as the full subcategory of all $M\in\operatorname{\mathsf{mod}}\varLambda$ filtered by quotient modules of $X$; that is, $M\in\operatorname{\mathsf{T}}(X)$ holds if and only if there exists a sequence
[…]
The subcategory $\operatorname{\mathsf{T}}(X)$ is the smallest torsion class containing $X$; here a torsion class means a full subcategory closed under taking quotient modules and extensions in $\operatorname{\mathsf{mod}}\varLambda$. This torsion class gives rise to the following torsion pair.
Lemma 2.7.
Let $X\in\operatorname{\mathsf{mod}}\varLambda$. Then we have a torsion pair $(\operatorname{\mathsf{T}}(X),X^{\perp})$ in $\operatorname{\mathsf{mod}}\varLambda$, that is, for each $M\in\operatorname{\mathsf{mod}}\varLambda$, there uniquely exists a short exact sequence
$0\to M_{1}\to M\to M_{2}\to 0\quad(M_{1}\in\operatorname{\mathsf{T}}(X),\ M_{2}\in X^{\perp}).$
The observation below is useful.
Proposition 2.8.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Assume that $X\in\operatorname{\mathsf{rep}}(Q,d)$ is $\theta$-stable for some $\theta\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}$. Then $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{T}}(X)=\mathcal{O}_{X}$.
Proof.
Clearly, $\mathcal{O}_{X}\subset\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{T}}(X)$ holds.
Let $M\in\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{T}}(X)$. To show $M\in\mathcal{O}_{X}$, assume the contrary $M\notin\mathcal{O}_{X}$. Since $M\in\operatorname{\mathsf{T}}(X)$, there exists a sequence
[…]
By the assumption $M\notin\mathcal{O}_{X}$, we get $X\not\simeq M$. Moreover $\operatorname{\underline{\mathsf{dim}}}X=d=\operatorname{\underline{\mathsf{dim}}}M$. Therefore we have $m\geq 2$ and each $M_{i}/M_{i-1}$ admits a surjection $X\to M_{i}/M_{i-1}$ which is not an isomorphism. Then $\theta(M_{i}/M_{i-1})>0$ holds for each $i$, because $X$ is $\theta$-stable. This and the filtration above imply $\theta(M)=\sum_{i=1}^{m}\theta(M_{i}/M_{i-1})>0$. Then since $\operatorname{\underline{\mathsf{dim}}}X=d=\operatorname{\underline{\mathsf{dim}}}M$, we get $\theta(X)=0$, but this contradicts that $X$ is $\theta$-stable. Thus $M\in\mathcal{O}_{X}$ holds. Therefore $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{T}}(X)\subset\mathcal{O}_{X}$ is true. ∎
This gives the next property.
Lemma 2.9.
Let $M\in\operatorname{\mathsf{mod}}\varLambda$. If $d:=\operatorname{\underline{\mathsf{dim}}}M\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ is a Schur root and $\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(M,M)\neq 0$, then there exists some $X\in\operatorname{\mathsf{mod}}\varLambda$ such that $\operatorname{\underline{\mathsf{dim}}}X=d$ and $M\notin\operatorname{\mathsf{T}}(X)$.
Proof.
Since $\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(M,M)\neq 0$, Lemma 2.6 gives that $\operatorname{\mathsf{rep}}(Q,d)\setminus\overline{\mathcal{O}_{M}}$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)$.
[…]
Thus there exists $X\in\operatorname{\mathsf{rep}}(Q,d)$ such that $X\notin\overline{\mathcal{O}_{M}}$ and that $X$ is $\theta$-stable.
It remains to show $M\notin\operatorname{\mathsf{T}}(X)$. Assume $M\in\operatorname{\mathsf{T}}(X)$. Then since $M\in\operatorname{\mathsf{rep}}(Q,d)$, Proposition 2.8 gives $M\in\mathcal{O}_{X}$. This is a contradiction, because $X\notin\overline{\mathcal{O}_{M}}\supset\mathcal{O}_{M}$. Thus we get $M\notin\operatorname{\mathsf{T}}(X)$. ∎
Then we have the following result.
Theorem 2.10.
Let $\varLambda=KQ$. Assume that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional, and set $d:=\operatorname{\underline{\mathsf{dim}}}B$. Then we have $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}B}\neq\emptyset$ and $\operatorname{\mathsf{rep}}(Q,d)\cap B^{\perp}\neq\emptyset$.
Proof.
We only prove $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}B}\neq\emptyset$. The other equality is similarly shown.
Since $B\in\operatorname{\mathsf{brick}}\varLambda$, its dimension vector $d$ is a Schur root. Then since $\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(B,B)\neq 0$, Lemma 2.9 allows us to take $X\in\operatorname{\mathsf{rep}}(Q,d)$ such that $B\notin\operatorname{\mathsf{T}}(X)$. Now Lemma 2.7 gives a short exact sequence
$0\to M_{1}\to B\to M_{2}\to 0\quad(M_{1}\in\operatorname{\mathsf{T}}(X),\ M_{2}\in X^{\perp}).$
We show that $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}M_{i}}\neq\emptyset$ for $i=1,2$. If this holds, then these sets are open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.2, so we get $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}B}$ as desired.
Since $B\notin\operatorname{\mathsf{T}}(X)$, we have $M_{1}\subsetneq B$. This and $B\in\operatorname{\mathsf{brick}}\varLambda$ imply $B\in{{}^{\perp}M_{1}}$; hence $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}M_{1}}$ is nonempty.
On the other hand, since $M_{2}\in X^{\perp}$ and $X\in\operatorname{\mathsf{rep}}(Q,d)$, we obtain $X\in\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}M_{2}}$; hence $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}M_{2}}$ is nonempty.
Thus we have the assertion. ∎
Now we can prove Theorem 1.3.
Proof.
Since $B$ is a brick which is not exceptional, both $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}B}$ and $\operatorname{\mathsf{rep}}(Q,d)\cap B^{\perp}$ are nonempty by Theorem 2.10. Then these sets are open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.2.
Moreover $B\in{{}^{\perp}M}\cap N^{\perp}$ and $B\in\operatorname{\mathsf{brick}}\varLambda$ implies that $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}M}$, $\operatorname{\mathsf{rep}}(Q,d)\cap N^{\perp}$ and $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ are nonempty. These sets are open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.2 again.
Therefore $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ is nonempty. Take $B^{\prime}$ from this set. Then we have the assertion. ∎
As in Section 1, Theorem 1.3 implies Theorem 1.1 and Corollary 1.2. ∎
We end this section with the following observations. Here a Schur root $d$ is said to be real if there exists some exceptional module $X\in\operatorname{\mathsf{rep}}(Q,d)$.
Remark 2.11.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ be a Schur root which is not real. As a subset of the direct product $\operatorname{\mathsf{rep}}(Q,d)\times\operatorname{\mathsf{rep}}(Q,d)$, we consider
$\mathcal{U}:=\{(X,Y)\in\operatorname{\mathsf{rep}}(Q,d)\times\operatorname{\mathsf{rep}}(Q,d)\mid\operatorname{\mathsf{Hom}}_{\varLambda}(X,Y)=0\}.$
By [S, Theorem 3.5] (cf. [MP, Proposition 3.4]), $\mathcal{U}$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)\times\operatorname{\mathsf{rep}}(Q,d)$; hence $\pi_{1}(\mathcal{U})$ and $\pi_{2}(\mathcal{U})$ are open dense subsets of $\operatorname{\mathsf{rep}}(Q,d)$, where $\pi_{1}$ and $\pi_{2}$ are the canonical projections $\operatorname{\mathsf{rep}}(Q,d)\times\operatorname{\mathsf{rep}}(Q,d)\to\operatorname{\mathsf{rep}}(Q,d)$. However, it is nontrivial that $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda\subset\pi_{1}(\mathcal{U})\cap\pi_{2}(\mathcal{U})$.
This inclusion follows from Theorem 2.10.
Corollary 2.12.
Let $\varLambda=KQ$, and $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ be a Schur root which is not real. Under the setting of Remark 2.11, we have
$\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda\subset\pi_{1}(\mathcal{U})\cap\pi_{2}(\mathcal{U}).$
Proof.
Let $B\in\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$. Since $d$ is a Schur root which is not real, $B$ is a brick which is not exceptional. By Theorem 2.10, there exists $X\in\operatorname{\mathsf{rep}}(Q,d)$ such that $\operatorname{\mathsf{Hom}}_{\varLambda}(B,X)=0$. Then $(B,X)\in\mathcal{U}$ holds, so we get $B\in\pi_{1}(\mathcal{U})$. Similarly, $B\in\pi_{2}(\mathcal{U})$ follows. ∎
3. Another proof of a weaker version of Theorem 1.3 via semi-invariants
In this section, we record our first proof of Corollary 1.2, since it is an interesting one relying on the well-established theory of semi-invariants [Ki, SV, DW] and presentation spaces [DF]. We aim to show the following weaker version of Theorem 1.3. The only difference is the existence of “$\mathbb{Z}_{\geq 1}$”.
Theorem 3.1.
Let $\varLambda=KQ$ with $Q$ a finite acyclic quiver, and $M,N\in\operatorname{\mathsf{mod}}\varLambda$. Assume that $B\in{{}^{\perp}M}\cap N^{\perp}$ and that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional. Then there exists some $B^{\prime}\in\operatorname{\mathsf{brick}}\varLambda$ such that
$B^{\prime}\in{{}^{\perp}(M\oplus B)}\cap(N\oplus B)^{\perp},\quad\operatorname{\underline{\mathsf{dim}}}B^{\prime}\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B.$
3.1. Preliminaries
In this subsection, we recall some additional results on quiver representations.
Lemma 3.2.
[Ka2, Proposition 1 (a)] Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then the following conditions are equivalent.
  1. (a)
    The element $d$ is a Schur root.
  2. (b)
    The set $\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid\text{$X$ is indecomposable}\}$ is a dense subset of $\operatorname{\mathsf{rep}}(Q,d)$.
By using Schur roots, we can consider canonical decompositions of dimension vectors.
Definition-Proposition 3.3.
[Ka1, Subsection 2.8] Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then there exist Schur roots $d_{1},\ldots,d_{m}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ such that the following set contains an open dense subset of $\operatorname{\mathsf{rep}}(Q,d)$:
$\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid\text{$X\simeq\bigoplus_{i=1}^{m}X_{i}$ with $\operatorname{\underline{\mathsf{dim}}}X_{i}=d_{i}$ and $X_{i}$ is indecomposable for each $i$}\}.$
Moreover such $d_{1},d_{2}\ldots,d_{m}$ are unique up to reordering. In this case, we write $d=\bigoplus_{i=1}^{m}d_{i}$, and call it the canonical decomposition of $d$.
The following trichotomy of Schur roots will play an important role. Here we used the characterization [Ka2, Proposition 3] of canonical decompositions.
Definition 3.4.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ be a Schur root.
  1. (1)
    We say that $d$ is real if there exists some exceptional module $X\in\operatorname{\mathsf{rep}}(Q,d)$.
  2. (2)
    We say that $d$ is tame if $d$ is not real but the canonical decomposition of $2d$ is $d\oplus d$.
  3. (3)
    We say that $d$ is wild if $d$ is neither real nor tame.
Let $d$ be a Schur root. If $d$ is real, then there uniquely exists a brick $B\in\operatorname{\mathsf{rep}}(Q,d)$ up to isomorphisms by [Ka1, Theorem 1], and $B$ is exceptional. Thus if there exists a brick $B\in\operatorname{\mathsf{rep}}(Q,d)$ which is not exceptional, then $d$ is tame or wild.
Canonical decompositions of multiples of Schur roots are given as follows.
Proposition 3.5.
[S, Theorem 3.8] Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ be a Schur root and $l\in\mathbb{Z}_{\geq 1}$. Then the canonical decomposition of $ld$ is $d^{\oplus l}$ if $d$ is real or tame, and $ld$ if $d$ is wild.
Proof.
The original proof of this result relies on [S, Theorem 3.3], so we need Crawley-Boevey’s correction [C2] also here, by the same reason we wrote in the proof of Proposition 2.5. Then the proofs of [S, Theorems 3.5, 3.7, 3.8] proceed. ∎
3.2. Proof of Theorem 3.1
As in the previous section, we use the Grothendieck group $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ and its dual abelian group $K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}=\operatorname{\mathsf{Hom}}_{\mathbb{Z}}(K_{0}(\operatorname{\mathsf{mod}}\varLambda),\mathbb{Z})$.
Moreover we identify $K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}$ with the (split) Grothendieck group $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ of the category $\operatorname{\mathsf{proj}}\varLambda$ of finitely generated projective modules. For each $i\in Q_{0}$, we write $P_{i}$ for the projective cover of the simple module $S_{i}$. Then $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ has the isoclasses of indecomposable projective modules $[P_{i}]$ for $i\in Q_{0}$ as a $\mathbb{Z}$-basis, so again we have $K_{0}(\operatorname{\mathsf{proj}}\varLambda)=\bigoplus_{i\in Q_{0}}\mathbb{Z}[P_{i}]\simeq\mathbb{Z}^{Q_{0}}$, where a bracket [][$\quad$][ ] is used only to denote an element of $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ in this section (elements of $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ will be expressed like $\operatorname{\underline{\mathsf{dim}}}M$ or $\operatorname{\underline{\mathsf{dim}}}M-\operatorname{\underline{\mathsf{dim}}}N$ by using dimension vectors of modules).
Via the Euler bilinear form !,?$\rangle\colon K_{0}(\operatorname{\mathsf{proj}}\varLambda)\times K_{0}(\operatorname{\mathsf{mod}}\varLambda)\to\mathbb{Z}$ given by $\langle[P_{i}],\operatorname{\underline{\mathsf{dim}}}S_{j}\rangle=\delta_{i,j}$, each element $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ induces a $\mathbb{Z}$-linear map $\theta:=\langle\theta,?\rangle\colon K_{0}(\operatorname{\mathsf{mod}}\varLambda)\to\mathbb{Z}$. In this way, we have $K_{0}(\operatorname{\mathsf{proj}}\varLambda)\simeq K_{0}(\operatorname{\mathsf{mod}}\varLambda)^{*}$. For a module $M\in\operatorname{\mathsf{mod}}\varLambda$, we set $\theta(M):=\theta(\operatorname{\underline{\mathsf{dim}}}M)\in\mathbb{Z}$ in short.
To each $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$, Derksen-Fei [DF] associated the presentation space $\operatorname{\mathsf{Hom}}(\theta)$.
Definition 3.6.
[DF, Definition 1.1] Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$.
  1. (1)
    We define $P^{0}_{\theta},P^{1}_{\theta}\in\operatorname{\mathsf{proj}}\varLambda$ up to isomorphisms so that $\theta=[P^{0}_{\theta}]-[P^{1}_{\theta}]$ and $\operatorname{\mathsf{add}}P^{0}_{\theta}\cap\operatorname{\mathsf{add}}P^{1}_{\theta}=\{0\}$.
  2. (2)
    We define an irreducible algebraic variety $\operatorname{\mathsf{Hom}}(\theta):=\operatorname{\mathsf{Hom}}_{\varLambda}(P_{1},P_{0})$, and call it the presentation space of $\theta$.
  3. (3)
[…]
Then the value $\theta(M)$ can be calculated as follows, where $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}\varLambda)$ is the bounded derived category over $\operatorname{\mathsf{mod}}\varLambda$. The following is well-known and elementary.
Lemma 3.7.
Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ and $f\in\operatorname{\mathsf{Hom}}(\theta)$. Take the maximum $P\in\operatorname{\mathsf{proj}}\varLambda$ such that $P[1]$ is a direct summand of $P_{f}$. Then for any $M\in\operatorname{\mathsf{mod}}\varLambda$, we have
$\theta(M)$ $=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}\varLambda)}(P_{f},M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}\varLambda)}(P_{f},M[1])$
$=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(\operatorname{\mathsf{Coker}}f,M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(P,M).$
We define a $\mathbb{Z}$-linear map
$\iota\colon K_{0}(\operatorname{\mathsf{proj}}\varLambda)\to K_{0}(\operatorname{\mathsf{mod}}\varLambda)$
so that $[P]\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ is sent to $\operatorname{\underline{\mathsf{dim}}}P\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ for every indecomposable projective module $P\in\operatorname{\mathsf{proj}}\varLambda$.
For any $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$, the set $\{f\in\operatorname{\mathsf{Hom}}(\theta)\mid\text{$f$ is injective}\}$ is an open subset of $\operatorname{\mathsf{Hom}}(\theta)$. Thus we have the following property.
Lemma 3.8.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}\simeq K_{0}(\operatorname{\mathsf{mod}}\varLambda)_{\geq 0}$, and set $\theta:=\iota^{-1}(d)$.
  1. (1)
    Let $N\in\operatorname{\mathsf{rep}}(Q,d)$. Taking its minimal projective presentation $0\to P_{1}\to P_{0}\to N\to 0$, we get $\iota^{-1}(d)=[P_{0}]-[P_{1}]$.
  2. (2)
    For any $l\in\mathbb{Z}_{\geq 1}$, the set $\{f\in\operatorname{\mathsf{Hom}}(l\theta)\mid\text{$f$ is injective}\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}(l\theta)$.
  3. (3)
    Let $N\in\operatorname{\mathsf{rep}}(Q,d)$ and $M\in\operatorname{\mathsf{mod}}\varLambda$. Then we have
    $\theta(M)=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(N,M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(N,M).$
Proof.
(1) By the short exact sequence $0\to P_{1}\to P_{0}\to N\to 0$, we have $d=\operatorname{\underline{\mathsf{dim}}}N=\operatorname{\underline{\mathsf{dim}}}P_{0}-\operatorname{\underline{\mathsf{dim}}}P_{1}=\iota([P_{0}])-\iota([P_{1}])$ in $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$. Since $\iota$ is an isomorphism, we have the assertion.
[…]
We use the following subcategory associated to each $\theta$.
Definition 3.9.
[BKT, Subsection 3.1] Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Then we set
¯$\overline{\mathcal{F}}_{\theta}$ :=$:=\{M\in\operatorname{\mathsf{mod}}\varLambda\mid\text{any submodule $L$ of $M$ satisfies $\theta(L)\leq 0$}\}.$
The subcategory ¯$\overline{\mathcal{F}}_{\theta}$ is a torsion-free class in $\operatorname{\mathsf{mod}}\varLambda$, that is, ¯$\overline{\mathcal{F}}_{\theta}$ is closed under taking extensions and submodules. However, we do not use this property directly. We have the following observation.
Lemma 3.10.
Assume that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional. Set $\theta:=\iota^{-1}(\operatorname{\underline{\mathsf{dim}}}B)\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Then we have $B\in\overline{\mathcal{F}}_{\theta}$.
Proof.
Since $B$ is a brick, we get $\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(B,B)=1$. Moreover, since $B$ is not exceptional, we obtain $\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(B,B)\geq 1$. Then Lemma 3.8 (3) gives $\theta(B)\leq 0$. For any proper submodule $L$ of $B$, we have $\operatorname{\mathsf{Hom}}_{\varLambda}(B,L)=0$, since $B$ is a brick. Then $\theta(L)\leq 0$ follows again from Lemma 3.8 (3). Thus we get $B\in\overline{\mathcal{F}}_{\theta}$. ∎
The following property from semi-invariants is important.
Proposition 3.11.
[F, Theorem 3.6] [AI, Theorem 4.3] [Gar, Proposition 4.10] Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ and $M\in\overline{\mathcal{F}}_{\theta}$. Then there exists some $l\in\mathbb{Z}_{\geq 1}$ such that $\{f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)\mid\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,M)=0\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$.
Lemma 3.10 and Proposition 3.11 give the next properties, a weaker version of Theorem 2.10.
Theorem 3.12.
Let $\varLambda=KQ$. Assume that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional.
  1. (1)
    There exists $X\in\operatorname{\mathsf{mod}}\varLambda$ satisfying $X\in{{}^{\perp}B}$ and $\operatorname{\underline{\mathsf{dim}}}X\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B$.
  2. (2)
    There exists $X\in\operatorname{\mathsf{mod}}\varLambda$ satisfying $X\in B^{\perp}$ and $\operatorname{\underline{\mathsf{dim}}}X\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B$.
Proof.
(1) We set $d:=\operatorname{\underline{\mathsf{dim}}}B$ and $\theta:=\iota^{-1}(d)\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Since $B$ is a brick which is not exceptional, Lemma 3.10 implies $B\in\overline{\mathcal{F}}_{\theta}$.
By Proposition 3.11, we can take $l\in\mathbb{Z}_{\geq 1}$ such that $\{f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)\mid\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,B)=0\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$. Moreover Lemma 3.8 (2) implies that $\{f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)\mid\text{$f$ is injective}\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$. Therefore we can take $f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$ which is injective and satisfies $\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,B)=0$.
Set $X:=\operatorname{\mathsf{Coker}}f$. Then $f$ is a minimal projective presentation of $X$, so we have $\iota^{-1}(\operatorname{\underline{\mathsf{dim}}}X)=l\theta$ in $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ by Lemma 3.8 (1). Thus we get $\operatorname{\underline{\mathsf{dim}}}X=\iota(l\theta)=ld=l\cdot\operatorname{\underline{\mathsf{dim}}}B$ in $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$. Moreover $\operatorname{\mathsf{Hom}}_{\varLambda}(X,B)=\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,B)=0$ are nothing but $X\in{{}^{\perp}B}$.
(2) is dually shown. ∎
Now we can prove Theorem 3.1.
Proof of Theorem 3.1.
Set $d:=\operatorname{\underline{\mathsf{dim}}}B$. Since $B$ is not exceptional, Theorem 3.12 allows us to take $X_{1},X_{2}\in\operatorname{\mathsf{mod}}\varLambda$ and $l_{1},l_{2}\in\mathbb{Z}_{\geq 1}$ with $X_{1}\in{{}^{\perp}B}$, $X_{2}\in{B^{\perp}}$ and $\operatorname{\underline{\mathsf{dim}}}X_{i}=l_{i}d$ ($i\in\{1,2\}$).
We first show that
$\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$
is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$.
Since $B\in{{}^{\perp}M}\cap{N^{\perp}}$, the module $B^{\oplus l_{1}l_{2}}$ belongs to both $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}M}$ and $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{N^{\perp}}$. Then by Lemma 2.2, these sets are open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$.
Moreover $X_{1}\in{{}^{\perp}B}$, $X_{2}\in{B^{\perp}}$ and $\operatorname{\underline{\mathsf{dim}}}X_{i}=l_{i}d$ ($i\in\{1,2\}$) imply $X_{1}^{\oplus l_{2}}\in{{}^{\perp}B}$, $X_{2}^{\oplus l_{1}}\in{B^{\perp}}$ and $\operatorname{\underline{\mathsf{dim}}}X_{1}^{\oplus l_{2}}=\operatorname{\underline{\mathsf{dim}}}X_{2}^{\oplus l_{1}}=l_{1}l_{2}d$. Thus the sets $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}B}$ and $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{B^{\perp}}$ are open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$ by Lemma 2.2 again.
Thus $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ is open dense.
Since $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional, $d$ is a Schur root which is tame or wild. Set $l:=1$ if $d$ is tame, and $l:=l_{1}l_{2}$ if $d$ is wild. It suffices to show that
$\mathcal{X}:=\operatorname{\mathsf{rep}}(Q,ld)\cap\operatorname{\mathsf{brick}}\varLambda\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$
is nonempty.
If $d$ is tame, then Proposition 3.5 and Definition-Proposition 3.3 imply that there exists an open dense subset $\mathcal{U}\subset\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$ such that
$\mathcal{U}\subset\{X\in\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\mid\text{$X\simeq\bigoplus_{i=1}^{l_{1}l_{2}}X_{i}$ with $\operatorname{\underline{\mathsf{dim}}}X_{i}=d$ and $X_{i}$ is indecomposable for each $i$}\}.$
Thus $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}\cap\mathcal{U}$ is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$. Take $X$ from this set, and take an indecomposable direct summand $Y$ of $X$. Then $Y\in{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ follows from $X\in{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$, and $\operatorname{\underline{\mathsf{dim}}}Y=d$ comes from $X\in\mathcal{U}$. Thus $Y$ belongs to $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$, so this set is open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.2. Moreover since $d$ is a Schur root, $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.2. Thus the set $\mathcal{X}\subset\operatorname{\mathsf{rep}}(Q,d)$ is open dense, and thus, nonempty.
If $d$ is wild, then $l_{1}l_{2}d$ is a Schur root by Proposition 3.5, so $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$. Since $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$, so is $\mathcal{X}$ as desired. ∎
Theorem 3.1 implies the following weaker version of Theorem 1.1.
Theorem 3.13.
Let $\varLambda=KQ$, and $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ be a finite semibrick. Assume that a brick $B\in\mathcal{S}$ is not exceptional. Then the following statements hold.
  1. (1)
    There exists some $B^{\prime}\in\operatorname{\mathsf{brick}}\varLambda\setminus\{B\}$ satisfying
    $\mathcal{S}\sqcup\{B^{\prime}\}\in\operatorname{\mathsf{sbrick}}\varLambda,\quad\operatorname{\underline{\mathsf{dim}}}B^{\prime}\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B.$
  2. (2)
    There exists an infinite semibrick $\mathcal{S}^{\prime}\in\operatorname{\mathsf{sbrick}}A$ such that
    $\mathcal{S}\subset\mathcal{S}^{\prime},\quad\text{$\operatorname{\underline{\mathsf{dim}}}B^{\prime}\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B$ for each $B^{\prime}\in\mathcal{S}^{\prime}\setminus\mathcal{S}$}.$
Then Corollary 1.2 follows from Theorem 3.13. ∎
Acknowlegments
The author deeply thanks Yuki Hirano for asking him whether Corollary 1.2 related to [HKO, Theorem 3.9] holds at MSJ Autumn Meeting 2024 in Osaka and giving him useful comments. He is grateful also to William Crawley-Boevey, Osamu Iyama, Kaveh Mousavand and Charles Paquette for improving drafts. The author was supported by JSPS KAKENHI Grant Number JP23K12957.