### v3

# Maximal finite semibricks consist only of open bricks

Abstract.
A semibrick is a set of modules satisfying Schur’s Lemma, and it is said to be maximal if it is not properly contained in another semibrick. For any finite dimensional algebra $\varLambda$ over an algebracally closed field $K$, we prove that any maximal finite semibrick $\mathcal{S}$ consists only of open bricks $B$, that is, bricks whose orbit closures $\overline{\mathcal{O}_{B}}$ are irreducible components in the representation schemes.
1. Our results
In this short paper, we discuss maximal finite semibricks in the module category $\operatorname{\mathsf{mod}}\varLambda$ of a finite dimensional algebra $\varLambda$ over an algebraically closed field $K$. Bricks, semibricks and related notions are defined as follows:
  • A $\varLambda$-module $B\in\operatorname{\mathsf{mod}}\varLambda$ is called a brick if the endomorphism ring $\operatorname{\mathsf{End}}_{\varLambda}(B)$ is isomorphic to $K$. We write $\operatorname{\mathsf{brick}}\varLambda$ for the set of isoclasses of bricks in $\operatorname{\mathsf{mod}}\varLambda$.
  • [A] A subset $\mathcal{S}\subset\operatorname{\mathsf{brick}}\varLambda$ is called a semibrick if $\operatorname{\mathsf{Hom}}_{\varLambda}(B,B^{\prime})=0$ for any $B\neq B^{\prime}\in\mathcal{S}$. We write $\operatorname{\mathsf{sbrick}}\varLambda$ for the set of semibricks in $\operatorname{\mathsf{mod}}\varLambda$.
  • A semibrick $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ is said to be finite if $\mathcal{S}$ is a finite set.
  • A semibrick $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ is said to be maximal if there exists no semibrick $\mathcal{S}^{\prime}$ such that $\mathcal{S}\subsetneq\mathcal{S}^{\prime}$.
It is well-known that there exist a finite quiver $Q$ and an admissible ideal $I\subset KQ$ such that $\varLambda$ is Morita equivalent to $KQ/I$, so we may assume that $\varLambda=KQ/I$.
To each $d=(d_{i})_{i\in Q_{0}}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$, we associate a scheme $\operatorname{\mathsf{rep}}(\varLambda,d)$ called the representation scheme (see Definition 2.1), which can be seen as the set of all $\varLambda$-modules $M$ whose dimension vectors $\operatorname{\underline{\mathsf{dim}}}M$ are $d$. We write $\operatorname{\mathsf{Irr}}(\varLambda,d)$ for the set of irreducible components of $\operatorname{\mathsf{rep}}(\varLambda,d)$, and set $\operatorname{\mathsf{Irr}}(\varLambda)$ as the disjoint union of all $\operatorname{\mathsf{Irr}}(\varLambda,d)$ for $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$.
The group $\operatorname{\mathsf{GL}}(d)=\prod_{i\in Q_{0}}\operatorname{\mathsf{GL}}(d_{i})$ acts naturally on $\operatorname{\mathsf{rep}}(\varLambda,d)$, and for each $M\in\operatorname{\mathsf{rep}}(\varLambda,d)$, its orbit $\mathcal{O}_{M}$ coincides with the subset $\{X\in\operatorname{\mathsf{rep}}(\varLambda,d)\mid\text{$X\simeq M$ as $\varLambda$-modules}\}$. Then every irreducible component of $\operatorname{\mathsf{rep}}(\varLambda,d)$ is a union of orbits, since $\operatorname{\mathsf{GL}}(d)$ is connected.
For a brick $B$, we say that a brick $B$ is an open brick if $\mathcal{O}_{B}$ is open dense in some $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$; otherwise, $B$ is called a non-open brick. The following is the main result of this paper.
Theorem 1.1.
Let $\varLambda$ be a finite dimensional algebra over an algebracally closed field $K$. Assume that $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ is a maximal finite semibrick. Then every brick in $\mathcal{S}$ is an open brick.
Theorem 1.1 is a direct consequence of the following general result, which is an extension theorem of semibricks.
Theorem 1.2.
Let $\varLambda$ be a finite dimensional algebra over an algebracally closed field $K$, and $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ be a finite semibrick. If $B\in\mathcal{S}$ and $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$ satisfy $\overline{\mathcal{O}_{B}}\subsetneq\mathcal{Z}$, then there exists some brick $B^{\prime}$ satisfying
$\mathcal{S}\sqcup\{B^{\prime}\}\in\operatorname{\mathsf{sbrick}}\varLambda,\quad B^{\prime}\in\mathcal{Z}.$
Applying Theorem 1.2 repeatedly, we get the following.
Corollary 1.3.
Let $\varLambda$ be a finite dimensional algebra over an algebracally closed field $K$, and $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ be a finite semibrick. Let $\mathcal{Z}_{1},\ldots,\mathcal{Z}_{m}\in\operatorname{\mathsf{Irr}}(\varLambda)$ with $m\geq 1$ satisfy that, for each $i\in\{1,\ldots,m\}$, there exists a non-open brick $B\in\mathcal{S}$ such that $B\in\mathcal{Z}_{i}$. Then there exist infinite semibricks $\mathcal{S}_{1},\ldots,\mathcal{S}_{m}$ satisfying
$\mathcal{S}\sqcup\mathcal{S}_{1}\sqcup\cdots\sqcup\mathcal{S}_{m}\in\operatorname{\mathsf{sbrick}}\varLambda,\quad\mathcal{S}_{i}\subset\mathcal{Z}_{i}\quad(i\in\{1,\ldots,m\}).$
The next section is devoted to showing Theorem 1.2.
In Section 3, we prove a weaker version of Proposition 2.6 for path algebras (Proposition 3.1) by using semi-invariants as a record, which was our first proof of Theorem 1.1.
2. Proofs of main results
2.1. Preliminaries
Throughout this paper, $K$ is an algebraically closed field, and $\varLambda=KQ/I$ is a finite dimensional algebra over $K$ with $Q=(Q_{0},Q_{1})$ a finite quiver and $I$ an admissible ideal of $KQ$. All finite dimensional algebras over $K$ are described in this way, up to Morita equivalences.
For each arrow $\alpha\colon i\to j$ in $Q$, we set $s(\alpha):=i$ and $t(\alpha):=j$, so $s(\alpha)$ and $t(\alpha)$ are the source and the target of $\alpha$. The composite of two arrows $\alpha\colon i\to j$ and $\beta\colon j\to k$ are denoted by $\alpha\beta$.
For any module $M\in\operatorname{\mathsf{mod}}\varLambda$, the element $\operatorname{\underline{\mathsf{dim}}}M:=(\operatorname{\mathsf{dim}}_{K}Me_{i})_{i\in Q_{0}}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ is called the dimension vector of $M$, where $e_{i}\in\varLambda$ is the idempotent corresponding to each $i\in Q_{0}$.
Now we recall the definition and some properties of representation schemes (also known as representation varieties) necessary in this paper. See [C3] for details.
Definition 2.1.
Let $d=(d_{i})_{i\in Q_{0}}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then we define the representation schemes $\operatorname{\mathsf{rep}}(Q,d)$ and $\operatorname{\mathsf{rep}}(\varLambda,d)$ of $Q$ and $\varLambda$ for $d$ by
$\operatorname{\mathsf{rep}}(Q,d)$ $:=\prod_{\alpha\in Q_{1}}\operatorname{\mathsf{Hom}}_{K}(K^{d_{s(\alpha)}},K^{d_{t(\alpha)}}),$
$\operatorname{\mathsf{rep}}(\varLambda,d)$ $:=\left\{(\varphi_{\alpha})_{\alpha\in Q_{1}}\in\operatorname{\mathsf{rep}}(Q,d)\mid\text{$(\varphi_{\alpha})_{\alpha\in Q_{1}}$ satisfies the relations corresponding to $I$}\right\}.$
Both are algebraic schemes equipped with the Zariski topology. We set $\operatorname{\mathsf{Irr}}(\varLambda,d)$ as the set of all irreducible components of $\operatorname{\mathsf{rep}}(\varLambda,d)$, and $\operatorname{\mathsf{Irr}}(\varLambda)$ as the disjoint union $\bigsqcup_{d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}}\operatorname{\mathsf{rep}}(\varLambda,d)$.
For $d=(d_{i})_{i\in Q_{0}}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$, the representation scheme $\operatorname{\mathsf{rep}}(\varLambda,d)$ admits the action of $\operatorname{\mathsf{GL}}(d):=\prod_{i\in Q_{0}}\operatorname{\mathsf{GL}}(d_{i})$ given by
$g\cdot M:=(g_{t(\alpha)}\varphi_{\alpha}g_{s(\alpha)}^{-1})_{\alpha\in Q_{1}}$
for any $M=(\varphi_{\alpha})_{\alpha\in Q_{1}}\in\operatorname{\mathsf{rep}}(\varLambda,d)$ and $g=(g_{i})_{i\in Q_{0}}$.
We set $\mathcal{O}_{M}$ as the orbit of each $M\in\operatorname{\mathsf{rep}}(\varLambda,d)$ with respect to $\operatorname{\mathsf{GL}}(d)$. Then we have
$\mathcal{O}_{M}=\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid\text{$X\simeq M$ as $\varLambda$-modules}\}.$
Since $\operatorname{\mathsf{GL}}(d)$ is a connected algebraic group, each orbit $\mathcal{O}_{M}$ is irreducible, so every irreducible component of $\operatorname{\mathsf{rep}}(\varLambda,d)$ is a union of orbits. In other words, if $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$ and $M\in\mathcal{Z}$, then the orbit $\mathcal{O}_{M}$ and its closure $\overline{\mathcal{O}_{M}}$ are contained in $\mathcal{Z}$. Moreover $\mathcal{O}_{M}$ is locally closed in $\operatorname{\mathsf{rep}}(\varLambda,d)$, so $\mathcal{O}_{M}$ is an open dense subset of $\overline{\mathcal{O}_{M}}$.
The next notion is the point of our results and arguments.
Definition 2.2.
Let $B\in\operatorname{\mathsf{brick}}\varLambda$. Then we say that $B$ is an open brick if $\mathcal{O}_{B}$ is an open dense subset of some $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$; or equivalently, $\overline{\mathcal{O}_{B}}\in\operatorname{\mathsf{Irr}}(\varLambda)$ holds. Otherwise, $B$ is called a non-open brick.
In particular, if $B\in\operatorname{\mathsf{brick}}\varLambda$ satisfies $B\in\mathcal{Z}_{1}\cap\mathcal{Z}_{2}$ for some distinct $\mathcal{Z}_{1},\mathcal{Z}_{2}\in\operatorname{\mathsf{Irr}}(\varLambda)$, then $B$ cannot be an open brick, because its orbit $\mathcal{O}_{B}$ is contained in $\mathcal{Z}_{1}\cap\mathcal{Z}_{2}$, which is a proper closed subset of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$.
Typical examples of open bricks are labeling bricks between two functorially finite torsion classes in $\operatorname{\mathsf{mod}}\varLambda$ [MP1, Theorem 6.1]. However, we do not use this fact in this paper.
2.2. Proof of Theorem 1.2
For any module $M\in\operatorname{\mathsf{mod}}\varLambda$, we set two subcategories
${{}^{\perp}M}:=\{X\in\operatorname{\mathsf{mod}}\varLambda\mid\operatorname{\mathsf{Hom}}_{\varLambda}(X,M)=0\},\quad M^{\perp}:=\{X\in\operatorname{\mathsf{mod}}\varLambda\mid\operatorname{\mathsf{Hom}}_{\varLambda}(M,X)=0\}.$
From upper semicontinuity, we get the following properties.
Lemma 2.3.
Let $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$, $M\in\operatorname{\mathsf{mod}}\varLambda$ and $t\in\mathbb{Z}_{\geq 0}$. Then the following sets are open in $\mathcal{Z}$:
$\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ $:=\{X\in\mathcal{Z}\mid X\in\operatorname{\mathsf{brick}}\varLambda\},$
$\mathcal{Z}_{M,\leq t}$ $:=\{X\in\mathcal{Z}\mid\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(X,M)\leq t\},$ $\mathcal{Z}\cap{{}^{\perp}M}$ $:=\{X\in\mathcal{Z}\mid X\in{{}^{\perp}M}\}=\mathcal{Z}_{M,\leq 0},$
$\mathcal{Z}^{M,\leq t}$ $:=\{X\in\mathcal{Z}\mid\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(M,X)\leq t\},$ $\mathcal{Z}\cap M^{\perp}$ $:=\{X\in\mathcal{Z}\mid X\in{M^{\perp}}\}=\mathcal{Z}^{M,\leq 0}.$
Thus $\mathcal{Z}_{M,t}:=\mathcal{Z}_{M,\leq t}\setminus\mathcal{Z}_{M,\leq t-1}$ and $\mathcal{Z}^{M,t}:=\mathcal{Z}^{M,\leq t}\setminus\mathcal{Z}^{M,\leq t-1}$ are locally closed subsets of $\mathcal{Z}$, where we set $\mathcal{Z}_{M,-1}=\mathcal{Z}^{M,-1}:=\emptyset$.
Let $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda,d)$ and $M\in\operatorname{\mathsf{rep}}(\varLambda,c)$ with $d,c\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. For every $X\in\mathcal{Z}$, we regard $\operatorname{\mathsf{Hom}}_{\varLambda}(X,M)$ as a subset of $\operatorname{\mathsf{Hom}}_{K}(K^{d},K^{c}):=\prod_{i\in Q_{0}}\operatorname{\mathsf{Hom}}_{K}(K^{d_{i}},K^{c_{i}})$. For any subset $\mathcal{A}\subset\mathcal{Z}$, we set
$\mathcal{Y}(\mathcal{A},M)$ $:=\{(X,f)\in\mathcal{Z}\times\operatorname{\mathsf{Hom}}_{K}(K^{d},K^{c})\mid X\in\mathcal{A},\ f\in\operatorname{\mathsf{Hom}}_{\varLambda}(X,M)\}.$
Then we have the following.
Lemma 2.4.
[B, Lemma 2.1] Let $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda,d)$, $M\in\operatorname{\mathsf{rep}}(\varLambda,c)$ with $d,c\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then for each $t\in\mathbb{Z}_{\geq 0}$, the set $\mathcal{Y}(\mathcal{Z}_{M,t},M)$ is a locally closed subset of $\mathcal{Z}\times\operatorname{\mathsf{Hom}}_{K}(K^{d},K^{c})$, and the natural projection
$\mathcal{Y}(\mathcal{Z}_{M,t},M)\to\mathcal{Z}_{M,t}$
is a vector bundle, and hence, an open map.
Proof.
The proof of [B, Lemma 2.1] is valid also in this situation. ∎
Then we obtain the next property.
Proposition 2.5.
Let $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$ and $B\in\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$. Then the following conditions are equivalent.
  1. (a)
    The set $\mathcal{Z}\cap{{}^{\perp}B}$ is open dense in $\mathcal{Z}$.
  2. (b)
    The set $\mathcal{Z}\cap{B^{\perp}}$ is open dense in $\mathcal{Z}$.
  3. (c)
    The orbit $\mathcal{O}_{B}$ is not open dense in $\mathcal{Z}$.
Proof.
[…]
Let (a) hold. Then $\mathcal{O}_{B}$ cannot be open dense, because otherwise, $(\mathcal{Z}\cap{{}^{\perp}B})\cap\mathcal{O}_{B}$ becomes open dense in $\mathcal{Z}$ by (a), which implies the contradiction $B\in{{}^{\perp}B}$. Thus we get (c).
Conversely, assume that (a) does not hold. Then $\mathcal{Z}\cap{{}^{\perp}B}$ is empty by Lemma 2.3. This and $B\in\operatorname{\mathsf{brick}}\varLambda$ imply $\mathcal{Z}_{B,\leq 1}=\mathcal{Z}_{B,1}\neq\emptyset$. Thus $\mathcal{U}:=\mathcal{Z}_{B,1}$ is open dense in $\mathcal{Z}$ by Lemma 2.3 again. Then Lemma 2.4 implies that the natural projection
$p\colon\mathcal{Y}(\mathcal{U},B)\to\mathcal{U}$
is an open map.
Define $\mathcal{V}:=\{(X,f)\in\mathcal{Y}(\mathcal{U},B)\mid\text{$f$ is an isomorphism}\}$. Then $\mathcal{V}$ is open in $\mathcal{Y}(\mathcal{U},B)$, and it is nonempty by $(B,1_{B})\in\mathcal{V}$, where $1_{B}$ is the identity map. Thus $p(\mathcal{V})$ is open and nonempty in $\mathcal{U}$, so $p(\mathcal{V})$ is open dense in $\mathcal{Z}$. On the other hand, $p(\mathcal{V})=\mathcal{O}_{B}$ holds by definition. Thus $\mathcal{O}_{B}$ is open dense in $\mathcal{Z}$, so (c) does not hold as desired. ∎
Now we can prove the following property.
Proposition 2.6.
Let $\varLambda$ be a finite dimensional algebra over an algebracally closed field $K$, and $M,N\in\operatorname{\mathsf{mod}}\varLambda$. Assume that $B\in\operatorname{\mathsf{brick}}\varLambda$ and $\mathcal{Z}\subset\operatorname{\mathsf{Irr}}(\varLambda)$ satisfy $B\in{{}^{\perp}M}\cap N^{\perp}$ and $\overline{\mathcal{O}_{B}}\subsetneq\mathcal{Z}$. Then there exists some brick $B^{\prime}$ such that
$B^{\prime}\in\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}.$
Proof.
Since $B$ is a brick such that $\overline{\mathcal{O}_{M}}\subsetneq\mathcal{Z}$, both $\mathcal{Z}\cap{{}^{\perp}B}$ and $\mathcal{Z}\cap B^{\perp}$ are open dense by Proposition 2.5. Moreover $B\in{{}^{\perp}M}\cap N^{\perp}$ and $B\in\operatorname{\mathsf{brick}}\varLambda$ imply that $\mathcal{Z}\cap{{}^{\perp}M}$, $\mathcal{Z}\cap N^{\perp}$ and $\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ are nonempty. These sets are open dense in $\mathcal{Z}$ by Lemma 2.3. Therefore we get
$\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}\neq\emptyset.$
Taking $B^{\prime}$ from this set, we have the assertion. ∎
Then Theorem 1.2 follows.
Proof of Theorem 1.2.
This is obtained by setting $M$ and $N$ as $\bigoplus_{X\in\mathcal{S}\setminus\{B\}}X$ in Proposition 2.6, because $\mathcal{S}$ is finite. ∎
As in Section 1, Theorem 1.2 implies Theorem 1.1 and Corollary 1.3. ∎
We end this section with an observation on brick components. As in [GLFS], an irreducible component $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$ is called a brick component if $\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda\neq\emptyset$. In this case, $\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\mathcal{Z}$ by Lemma 2.3.
Proposition 2.7.
[MP3, Proposition 3.4] Let $\mathcal{Z}\in\operatorname{\mathsf{Irr}}(\varLambda)$ be a brick component. Then exactly one of the following holds.
  1. (a)
    There exists an open brick $B\in\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ such that $\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda=\mathcal{O}_{B}$; in other words, $B$ is the unique brick in $\mathcal{Z}$ up to isomorphisms.
  2. (b)
    The set $\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ is a union of infinitely many non-open bricks.
Thus open bricks are nothing but the bricks appearing in (a).
In [MP3], Proposition 2.7 is proved by using [MP3, Remark 3.3]; namely, every $B\in\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ satisfies $\operatorname{\mathsf{dim}}_{K}\mathcal{O}_{B}=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{GL}}(d)-1$. We give another proof of this proposition with Proposition 2.5.
Proof.
It suffices to show the following statement:
if there exists an open brick $B$ in $\mathcal{Z}$, then $\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda\subset\mathcal{O}_{B}$.
Let $B$ be an open brick in $\mathcal{Z}$. Assume that there exists $B^{\prime}\in\mathcal{Z}\cap\operatorname{\mathsf{brick}}\varLambda$ such that $B^{\prime}\notin\mathcal{O}_{B}$. Then since $\mathcal{O}_{B}$ is open dense in $\mathcal{Z}$, the orbit $\mathcal{O}_{B^{\prime}}\subset\mathcal{Z}$ cannot be open dense in $\mathcal{Z}$, so Proposition 2.5 implies that $\mathcal{Z}\cap{{}^{\perp}(B^{\prime})}$ is open dense in $\mathcal{Z}$. This set is a union of orbits, so must contain the open dense orbit $\mathcal{O}_{B}$. Thus we get $B\in{{}^{\perp}(B^{\prime})}$, which implies $B^{\prime}\in B^{\perp}$, so Lemma 2.3 gives $\mathcal{Z}\cap B^{\perp}$ is open dense in $\mathcal{Z}$. By Proposition 2.5 again, $\mathcal{O}_{B}$ is not open dense in $\mathcal{Z}$, a contradiction. ∎
3. Proposition 2.6 for path algebras
In this section, we record our first proof of a weaker version of Proposition 2.6 for path algebras, since it is an interesting one relying on the well-established theory of semi-invariants [Ki, SV, DW] and presentation spaces [DF]. Theorem 1.1 for path algebras follows also from this.
We assume that $\varLambda=KQ$ for a finite acyclic quiver $Q$ in the rest. We say that a brick $B\in\operatorname{\mathsf{brick}}\varLambda$ is exceptional if $\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(B,B)=0$ as in [C1]. Since $\varLambda$ is a path algebra, a brick $B\in\operatorname{\mathsf{brick}}\varLambda$ is an open brick if and only if $B$ is exceptional by Voigt’s Lemma [Gab, V].
We aim to show the following weaker version of Proposition 2.6.
Proposition 3.1.
Let $\varLambda=KQ$ with $Q$ a finite acyclic quiver, and $M,N\in\operatorname{\mathsf{mod}}\varLambda$. Assume that $B\in{{}^{\perp}M}\cap N^{\perp}$ and that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional. Then there exists some $B^{\prime}\in\operatorname{\mathsf{brick}}\varLambda$ such that
$B^{\prime}\in{{}^{\perp}(M\oplus B)}\cap(N\oplus B)^{\perp},\quad\operatorname{\underline{\mathsf{dim}}}B^{\prime}\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B.$
3.1. Preliminaries
In this subsection, we recall some additional properties on quiver representations.
Since $\varLambda=KQ$, the representation scheme $\operatorname{\mathsf{rep}}(\varLambda,d)$ for $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ is just $\operatorname{\mathsf{rep}}(Q,d)$. It is the direct product $\prod_{\alpha\in Q_{1}}\operatorname{\mathsf{Hom}}_{K}(K^{s(\alpha),t(\alpha)})$, so in particular, irreducible. Thus $\operatorname{\mathsf{Irr}}(\varLambda)$ is identified with $(\mathbb{Z}_{\geq 0})^{Q_{0}}$.
We define Schur roots as follows.
Definition 3.2.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then we call $d$ a Schur root if $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ is nonempty.
By Lemma 2.3, if $d$ is a Schur root, then $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)$.
Schur roots depend on the orientation of the quiver $Q$, but do not depend on the field $K$. This follows from [S, Proposition 6.1] together with the correction [C2] of [S, Proposition 3.3].
We recall the following basic property of path algebras.
Lemma 3.3.
[Ka2, Proposition 1 (a)] Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then the following conditions are equivalent.
  1. (a)
    The element $d$ is a Schur root.
  2. (b)
    The set $\{X\in\operatorname{\mathsf{rep}}(Q,d)\mid\text{$X$ is indecomposable}\}$ contains an open dense subset of $\operatorname{\mathsf{rep}}(Q,d)$.
By using Schur roots, we can consider canonical decompositions of dimension vectors.
Definition-Proposition 3.4.
[Ka1, Subsection 2.8 a)] Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Then there exist Schur roots $d_{1},\ldots,d_{m}\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ such that the following set contains an open dense subset of $\operatorname{\mathsf{rep}}(Q,d)$:
[…]
Moreover such $d_{1},d_{2}\ldots,d_{m}$ are unique up to reordering. In this case, we write $d=\bigoplus_{i=1}^{m}d_{i}$, and call it the canonical decomposition of $d$.
The following trichotomy of Schur roots will play an important role. Here we used the characterization [Ka2, Proposition 3] of canonical decompositions.
Definition 3.5.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ be a Schur root.
  1. (1)
    We say that $d$ is real if there exists some exceptional module $X\in\operatorname{\mathsf{rep}}(Q,d)$.
  2. (2)
    We say that $d$ is tame if $d$ is not real but the canonical decomposition of $2d$ is $d\oplus d$.
  3. (3)
    We say that $d$ is wild if $d$ is neither real nor tame.
Let $d$ be a Schur root. If $d$ is real, then there uniquely exists a brick $B\in\operatorname{\mathsf{rep}}(Q,d)$ up to isomorphisms by [Ka1, Theorem 1], and $B$ is exceptional. Thus if there exists a brick $B\in\operatorname{\mathsf{rep}}(Q,d)$ which is not exceptional, then $d$ is tame or wild.
Canonical decompositions of multiples of Schur roots are given as follows.
Proposition 3.6.
[S, Theorem 3.8] Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$ be a Schur root and $l\in\mathbb{Z}_{\geq 1}$. Then the canonical decomposition of $ld$ is $d^{\oplus l}$ if $d$ is real or tame, and $ld$ if $d$ is wild.
Proof.
The original proof of [S, Theorem 3.3] works in the case that the characteristic of $K$ is zero, but it still holds for any algebraically closed field thanks to Crawley-Boevey’s correction [C2]. Then the proofs of [S, Theorems 3.5, 3.7, 3.8] proceed. ∎
3.2. Proof of Proposition 3.1
We will use the Grothendieck group $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ here.
For each vertex $i\in Q_{0}$, we write $S_{i}$ for the simple $\varLambda$-module corresponding to $i$, and $S_{i}]$ for its isoclass. Then the Grothendieck group $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ is the free abelian group $\bigoplus_{i\in Q_{0}}\mathbb{Z}[S_{i}]$, so we get $K_{0}(\operatorname{\mathsf{mod}}\varLambda)\simeq\mathbb{Z}^{Q_{0}}$ as abelian groups. Under this isomorphism, for each module $M\in\operatorname{\mathsf{mod}}\varLambda$, the element $M]\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ corresponds to its dimension vector $\operatorname{\underline{\mathsf{dim}}}M\in(\mathbb{Z}_{\geq 0})^{Q_{0}}$. Thus in what follows, the dimension vector $\operatorname{\underline{\mathsf{dim}}}M$ is regarded as an element of $K_{0}(\operatorname{\mathsf{mod}}\varLambda)_{\geq 0}:=\sum_{i\in Q_{0}}\mathbb{Z}_{\geq 0}[S_{i}]$.
We also use the (split) Grothendieck group $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ of the category $\operatorname{\mathsf{proj}}\varLambda$ of finitely generated projective modules. For each $i\in Q_{0}$, we write $P_{i}$ for the projective cover of the simple module $S_{i}$. Then $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ has the isoclasses of indecomposable projective modules $P_{i}]$ for $i\in Q_{0}$ as a $\mathbb{Z}$-basis, so again we have $K_{0}(\operatorname{\mathsf{proj}}\varLambda)=\bigoplus_{i\in Q_{0}}\mathbb{Z}[P_{i}]\simeq\mathbb{Z}^{Q_{0}}$.
From now on, a bracket $\quad$] will be used only to denote an element of $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Elements of $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ will be expressed like $\operatorname{\underline{\mathsf{dim}}}M$ or $\operatorname{\underline{\mathsf{dim}}}M-\operatorname{\underline{\mathsf{dim}}}N$ by using dimension vectors of modules.
Via the Euler bilinear form $\langle!,?\rangle\colon K_{0}(\operatorname{\mathsf{proj}}\varLambda)\times K_{0}(\operatorname{\mathsf{mod}}\varLambda)\to\mathbb{Z}$ given by $\langle[P_{i}],\operatorname{\underline{\mathsf{dim}}}S_{j}\rangle=\delta_{i,j}$, each element $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ induces a $\mathbb{Z}$-linear map $\theta:=\langle\theta,?\rangle\colon K_{0}(\operatorname{\mathsf{mod}}\varLambda)\to\mathbb{Z}$. In this way, we have $K_{0}(\operatorname{\mathsf{proj}}\varLambda)\simeq\operatorname{\mathsf{Hom}}_{\mathbb{Z}}(K_{0}(\operatorname{\mathsf{mod}}\varLambda),\mathbb{Z})$. For a module $M\in\operatorname{\mathsf{mod}}\varLambda$, we set $\theta(M):=\theta(\operatorname{\underline{\mathsf{dim}}}M)\in\mathbb{Z}$ in short.
To each $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$, Derksen-Fei [DF] associated the presentation space $\operatorname{\mathsf{Hom}}(\theta)$.
Definition 3.7.
[DF, Definition 1.1] Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$.
  1. (1)
    We define $P^{0}_{\theta},P^{1}_{\theta}\in\operatorname{\mathsf{proj}}\varLambda$ up to isomorphisms so that $\theta=[P^{0}_{\theta}]-[P^{1}_{\theta}]$ and $\operatorname{\mathsf{add}}P^{0}_{\theta}\cap\operatorname{\mathsf{add}}P^{1}_{\theta}=\{0\}$.
  2. (2)
    We define an irreducible algebraic variety $\operatorname{\mathsf{Hom}}(\theta):=\operatorname{\mathsf{Hom}}_{\varLambda}(P_{1},P_{0})$, and call it the presentation space of $\theta$.
  3. (3)
    For each $f\in\operatorname{\mathsf{Hom}}(\theta)$, a 2-term complex $P_{f}$ is defined by $P_{f}:=(P_{1}^{\theta}\xrightarrow{f}P_{0}^{\theta})$ in the homotopy category $\mathsf{K}^{\mathrm{b}}(\operatorname{\mathsf{proj}}\varLambda)$ of bounded complexes over $\operatorname{\mathsf{proj}}\varLambda$.
Then the value $\theta(M)$ can be calculated as follows, where $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}\varLambda)$ is the bounded derived category over $\operatorname{\mathsf{mod}}\varLambda$. The following is well-known and elementary.
Lemma 3.8.
Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ and $f\in\operatorname{\mathsf{Hom}}(\theta)$. Take the maximum $P\in\operatorname{\mathsf{proj}}\varLambda$ such that $P[1]$ is a direct summand of $P_{f}$. Then for any $M\in\operatorname{\mathsf{mod}}\varLambda$, we have
$\theta(M)$ $=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}\varLambda)}(P_{f},M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}\varLambda)}(P_{f},M[1])$
$=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(\operatorname{\mathsf{Coker}}f,M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(P,M).$
We define a $\mathbb{Z}$-linear map
$\iota\colon K_{0}(\operatorname{\mathsf{proj}}\varLambda)\to K_{0}(\operatorname{\mathsf{mod}}\varLambda)$
so that $P]\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ is sent to $\operatorname{\underline{\mathsf{dim}}}P\in K_{0}(\operatorname{\mathsf{mod}}\varLambda)$ for every indecomposable projective module $P\in\operatorname{\mathsf{proj}}\varLambda$.
For any $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$, the set $\{f\in\operatorname{\mathsf{Hom}}(\theta)\mid\text{$f$ is injective}\}$ is an open subset of $\operatorname{\mathsf{Hom}}(\theta)$. Thus we have the following property.
Lemma 3.9.
Let $d\in(\mathbb{Z}_{\geq 0})^{Q_{0}}\simeq K_{0}(\operatorname{\mathsf{mod}}\varLambda)_{\geq 0}$, and set $\theta:=\iota^{-1}(d)$.
  1. (1)
    Let $N\in\operatorname{\mathsf{rep}}(Q,d)$. Taking its minimal projective presentation $0\to P_{1}\to P_{0}\to N\to 0$, we get $\iota^{-1}(d)=[P_{0}]-[P_{1}]$.
  2. (2)
    For any $l\in\mathbb{Z}_{\geq 1}$, the set $\{f\in\operatorname{\mathsf{Hom}}(l\theta)\mid\text{$f$ is injective}\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}(l\theta)$.
  3. (3)
    Let $N\in\operatorname{\mathsf{rep}}(Q,d)$ and $M\in\operatorname{\mathsf{mod}}\varLambda$. Then we have
    $\theta(M)=\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(N,M)-\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(N,M).$
Proof.
(1) By the short exact sequence $0\to P_{1}\to P_{0}\to N\to 0$, we have $d=\operatorname{\underline{\mathsf{dim}}}N=\operatorname{\underline{\mathsf{dim}}}P_{0}-\operatorname{\underline{\mathsf{dim}}}P_{1}=\iota([P_{0}])-\iota([P_{1}])$ in $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$. Since $\iota$ is an isomorphism, we have the assertion.
(2) By upper semicontinuity, we know that $\{f\in\operatorname{\mathsf{Hom}}(\theta)\mid\text{$f$ is injective}\}$ is open in $\operatorname{\mathsf{Hom}}(l\theta)$, so will show that it is nonempty. Take $N\in\operatorname{\mathsf{rep}}(Q,d)$ and its minimal presentation $0\to P_{1}\xrightarrow{f}P_{0}\to N\to 0$. Then $f^{\oplus l}$ is injective. Thus $\{f\in\operatorname{\mathsf{Hom}}(\theta)\mid\text{$f$ is injective}\}$ is nonempty; hence it is open dense in $\operatorname{\mathsf{Hom}}(l\theta)$.
(3) Take the minimal presentation $0\to P_{1}\xrightarrow{f}P_{0}\to N\to 0$. Then $f$ is injective, so Lemma 3.8 implies the assertion. ∎
We use the following subcategory associated to each $\theta$.
Definition 3.10.
[BKT, Subsection 3.1] Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Then we set
$\overline{\mathcal{F}}_{\theta}$ $:=\{M\in\operatorname{\mathsf{mod}}\varLambda\mid\text{any submodule $L$ of $M$ satisfies $\theta(L)\leq 0$}\}.$
The subcategory $\overline{\mathcal{F}}_{\theta}$ is a torsion-free class in $\operatorname{\mathsf{mod}}\varLambda$, that is, $\overline{\mathcal{F}}_{\theta}$ is closed under taking extensions and submodules. However, we do not use this property directly. We have the following observation.
Lemma 3.11.
Assume that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional. Set $\theta:=\iota^{-1}(\operatorname{\underline{\mathsf{dim}}}B)\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Then we have $B\in\overline{\mathcal{F}}_{\theta}$.
Proof.
Since $B$ is a brick, we get $\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Hom}}_{\varLambda}(B,B)=1$. Moreover, since $B$ is not exceptional, we obtain $\operatorname{\mathsf{dim}}_{K}\operatorname{\mathsf{Ext}}_{\varLambda}^{1}(B,B)\geq 1$. Then Lemma 3.9 (3) gives $\theta(B)\leq 0$. For any proper submodule $L$ of $B$, we have $\operatorname{\mathsf{Hom}}_{\varLambda}(B,L)=0$, since $B$ is a brick. Then $\theta(L)\leq 0$ follows again from Lemma 3.9 (3). Thus we get $B\in\overline{\mathcal{F}}_{\theta}$. ∎
The following property from semi-invariants is important.
Proposition 3.12.
[F, Theorem 3.6] [AI, Theorem 4.3] [Gar, Proposition 4.10] Let $\theta\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ and $M\in\overline{\mathcal{F}}_{\theta}$. Then there exists some $l\in\mathbb{Z}_{\geq 1}$ such that $\{f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)\mid\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,M)=0\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$.
Lemma 3.11 and Proposition 3.12 give the next properties.
Proposition 3.13.
Let $\varLambda=KQ$. Assume that $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional.
  1. (1)
    There exists $X\in\operatorname{\mathsf{mod}}\varLambda$ satisfying $X\in{{}^{\perp}B}$ and $\operatorname{\underline{\mathsf{dim}}}X\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B$.
  2. (2)
    There exists $X\in\operatorname{\mathsf{mod}}\varLambda$ satisfying $X\in B^{\perp}$ and $\operatorname{\underline{\mathsf{dim}}}X\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B$.
Proof.
(1) We set $d:=\operatorname{\underline{\mathsf{dim}}}B$ and $\theta:=\iota^{-1}(d)\in K_{0}(\operatorname{\mathsf{proj}}\varLambda)$. Since $B$ is a brick which is not exceptional, Lemma 3.11 implies $B\in\overline{\mathcal{F}}_{\theta}$.
By Proposition 3.12, we can take $l\in\mathbb{Z}_{\geq 1}$ such that $\{f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)\mid\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,B)=0\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$. Moreover Lemma 3.9 (2) implies that $\{f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)\mid\text{$f$ is injective}\}$ is an open dense subset of $\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$. Therefore we can take $f\in\operatorname{\mathsf{Hom}}_{\varLambda}(l\theta)$ which is injective and satisfies $\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,B)=0$.
Set $X:=\operatorname{\mathsf{Coker}}f$. Then $f$ is a minimal projective presentation of $X$, so we have $\iota^{-1}(\operatorname{\underline{\mathsf{dim}}}X)=l\theta$ in $K_{0}(\operatorname{\mathsf{proj}}\varLambda)$ by Lemma 3.9 (1). Thus we get $\operatorname{\underline{\mathsf{dim}}}X=\iota(l\theta)=ld=l\cdot\operatorname{\underline{\mathsf{dim}}}B$ in $K_{0}(\operatorname{\mathsf{mod}}\varLambda)$. Moreover $\operatorname{\mathsf{Hom}}_{\varLambda}(X,B)=\operatorname{\mathsf{Hom}}_{\varLambda}(\operatorname{\mathsf{Coker}}f,B)=0$ are nothing but $X\in{{}^{\perp}B}$.
(2) is dually shown. ∎
Now we can prove Proposition 3.1.
Proof of Proposition 3.1.
Set $d:=\operatorname{\underline{\mathsf{dim}}}B$. Since $B$ is not exceptional, Proposition 3.13 allows us to take $X_{1},X_{2}\in\operatorname{\mathsf{mod}}\varLambda$ and $l_{1},l_{2}\in\mathbb{Z}_{\geq 1}$ with $X_{1}\in{{}^{\perp}B}$, $X_{2}\in{B^{\perp}}$ and $\operatorname{\underline{\mathsf{dim}}}X_{i}=l_{i}d$ ($i\in\{1,2\}$).
We first show that
$\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$
is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$.
Since $B\in{{}^{\perp}M}\cap{N^{\perp}}$, the module $B^{\oplus l_{1}l_{2}}$ belongs to both $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}M}$ and $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{N^{\perp}}$. Then by Lemma 2.3, these sets are open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$.
Moreover $X_{1}\in{{}^{\perp}B}$, $X_{2}\in{B^{\perp}}$ and $\operatorname{\underline{\mathsf{dim}}}X_{i}=l_{i}d$ ($i\in\{1,2\}$) imply $X_{1}^{\oplus l_{2}}\in{{}^{\perp}B}$, $X_{2}^{\oplus l_{1}}\in{B^{\perp}}$ and $\operatorname{\underline{\mathsf{dim}}}X_{1}^{\oplus l_{2}}=\operatorname{\underline{\mathsf{dim}}}X_{2}^{\oplus l_{1}}=l_{1}l_{2}d$. Thus the sets $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}B}$ and $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{B^{\perp}}$ are open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$ by Lemma 2.3 again.
Thus $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ is open dense.
Since $B\in\operatorname{\mathsf{brick}}\varLambda$ is not exceptional, $d$ is a Schur root which is tame or wild. Set $l:=1$ if $d$ is tame, and $l:=l_{1}l_{2}$ if $d$ is wild. It suffices to show that
$\mathcal{X}:=\operatorname{\mathsf{rep}}(Q,ld)\cap\operatorname{\mathsf{brick}}\varLambda\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$
is nonempty.
If $d$ is tame, then Proposition 3.6 and Definition-Proposition 3.4 imply that there exists an open dense subset $\mathcal{U}\subset\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$ such that
[…]
Thus $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}\cap\mathcal{U}$ is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$. Take $X$ from this set, and take an indecomposable direct summand $Y$ of $X$. Then $Y\in{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ follows from $X\in{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$, and $\operatorname{\underline{\mathsf{dim}}}Y=d$ comes from $X\in\mathcal{U}$. Thus $Y$ belongs to $\operatorname{\mathsf{rep}}(Q,d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$, so this set is open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.3. Moreover since $d$ is a Schur root, $\operatorname{\mathsf{rep}}(Q,d)\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\operatorname{\mathsf{rep}}(Q,d)$ by Lemma 2.3. Thus the set $\mathcal{X}\subset\operatorname{\mathsf{rep}}(Q,d)$ is open dense, and thus, nonempty.
If $d$ is wild, then $l_{1}l_{2}d$ is a Schur root by Proposition 3.6, so $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap\operatorname{\mathsf{brick}}\varLambda$ is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$. Since $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)\cap{{}^{\perp}(M\oplus B)}\cap{(N\oplus B)^{\perp}}$ is open dense in $\operatorname{\mathsf{rep}}(Q,l_{1}l_{2}d)$, so is $\mathcal{X}$ as desired. ∎
Proposition 3.1 implies the following weaker version of Theorem 1.2 for path algebras.
Theorem 3.14.
Let $\varLambda=KQ$, and $\mathcal{S}\in\operatorname{\mathsf{sbrick}}\varLambda$ be a finite semibrick. Assume that a brick $B\in\mathcal{S}$ is not exceptional. Then there exists some brick $B^{\prime}$ satisfying
$\mathcal{S}\sqcup\{B^{\prime}\}\in\operatorname{\mathsf{sbrick}}\varLambda,\quad\operatorname{\underline{\mathsf{dim}}}B^{\prime}\in\mathbb{Z}_{\geq 1}\cdot\operatorname{\underline{\mathsf{dim}}}B.$
Then Theorem 1.1 for path algebras follows from Theorem 3.14. ∎
Acknowlegments
The author deeply thanks Yuki Hirano for asking him whether Theorem 1.1 for path algebras related to [HKO, Theorem 3.9] holds at MSJ Autumn Meeting 2024 in Osaka and giving him useful comments. He appreciates the helpful comments of an anonymous referee. He is grateful also to William Crawley-Boevey, Osamu Iyama, Kaveh Mousavand and Charles Paquette for improving drafts. The author was supported by JSPS KAKENHI Grant Number JP23K12957.